# Diffence: Fencing Membership Privacy With Diffusion Models

### v1

@@ 3.1 Prediction Gaps between Members and Non-members
Figure 3 plots histograms illustrating the differences in predictions for members and non-members on CIFAR-100 without defenses. It can be observed that there are significant disparities in the output distributions between members and non-members, especially in the parametric modeled confidence which we refer to as logit in this paper. Attackers can easily exploit these differences to distinguish between members and non-members.
3.2 Proposed Diffence
To solve the issues described above, we propose a novel method that leverages input reconstruction to eliminate the prediction disparity between members and non-members. Our core idea is that by reconstructing samples, the model, during classification, encounters samples that are distinct from those it has seen during the training phase, irrespective of whether they are from members or non-members, thereby reducing inconsistencies in predictions. Our method aims to reconstruct the finer details of samples without altering their semantic content. To this end, we employ diffusion models as our generative model, as they align well with our objectives. We provide more details below.
3.2.1 Sample Reconstruction
@@
Typically, images display a high response to low-frequency content and a notably weaker response to high-frequency content. This is because of the inherent smoothness of most images. The dominant low-frequency components encapsulate essential visual information, while the high-frequency components, associated with fine details and edges, are comparatively subdued.
@@
In our defense strategy, we denote the diffusion step applied to each sample as $T$. By selecting an appropriate value for $T$, we can preserve the essential semantic content of each image while altering its finer details. The impact of the diffusion step $T$ on the performance of our defense is further explored in Section 4.3.2. During the inference phase, the two-step reconstruction procedure reduces the gap between the prediction distributions of members and non-members.
3.2.2 Sample Selection
@@
Figure 4: Prediction distribution gap between members and non-members under our defense across the three scenarios. This figure displays the distribution of predicted logits on the CIFAR100 dataset using ResNet with our defense in the three scenarios.
We observed that the stochastic nature of sample generation can lead to a decline in sample quality sometimes when only one reconstructed sample is generated for each image, potentially resulting in decreased model test accuracy. To address this issue, we propose generating multiple reconstructed images for each original image. Then in alignment with the specific needs of our defense strategy, we carefully select the most appropriate samples to be utilized as the final input sample.
In our approach, for each original image $I_{i}$ requiring privacy protection, we generate $N$ reconstructed versions, denoted as $R_{i}=\{R_{i1},R_{i2},\ldots,R_{iN}\}$. The model $f$ outputs a prediction for each reconstructed image $R_{ij}$, i.e., $f(R_{ij})$. From these, we identify a subset of reconstructions whose predicted labels match the label of the original image $I_{i}$ forming a set of candidate reconstructions, denoted as $C$. We then select the most appropriate reconstructed image from the candidates. Selecting a sample from this candidate set ensures that the original prediction label of the sample remains unchanged, guaranteeing that our method does not impact the model’s accuracy.
@@
Here, $P_{i}$ represents the final prediction for the original image $I_{i}$, derived from the model’s prediction for the selected reconstruction. The selection function, denoted as $\text{select}(\cdot)$, chooses the optimal reconstruction from $R_{i}$ based on a criterion that evaluates the predictions of all $N$ reconstructions. This criterion is aligned with the defense’s objectives and may take into account the defender’s prior knowledge about the prediction distributions of members and non-members. We categorize our defense into three scenarios based on the information available to the defender and have accordingly designed distinct sample selection strategies for each.
Scenario 1. In this scenario, we assume that the defender has access to subsets of both member and non-member data. The defender initially generates $N$ reconstructed samples for each data point and then plots the prediction distributions for all these reconstructions. Through the analysis of these distribution plots, the defender can identify an optimal interval where the predictions should ideally fall to maximize the reduction of the prediction distribution gap between members and non-members. Specifically, a grid search is conducted over the overlapping regions of member and non-member prediction distribution $[min(logit_{mem}),max(logit_{nonmem})]$ to determine this interval. The chosen interval aims to minimize the Jensen-Shannon (JS) divergence [31] between the prediction distributions of members and non-members within the selected range. This approach is based on the observation that a lower JS divergence between member and non-member prediction distributions typically indicates reduced membership privacy leakage [14]. This correlation has been confirmed by our experiments, as detailed in Section 4.3.1.
Defenders require only a small subset of samples to select the optimal interval. After the interval is determined, for each sample requiring protection, the defender randomly selects one sample from the candidate set $C$ that falls within this interval based on the prediction. If none of the candidates are within the interval, the closest one is chosen. Another possible approach involves continuously generating reconstructions until the sample falls within the defined interval. However, we found this method to be inefficient. More details on this are discussed in Section 4.3.1.
Figure 4 provides an example of our defense strategy as applied to the CIFAR-100 dataset. It illustrates the prediction distribution of reconstructed samples for 1000 data points which is used to select the optimal interval. The figure also showcases the distribution of prediction logits for both members and non-members after the selection of the optimal interval using the steps outlined in our defense strategy. This visualization clearly demonstrates the effectiveness of our approach in aligning the prediction distributions of member and non-member data, thereby minimizing the potential for privacy leakage.
@@
Scenario 3. In this scenario, the defender is unaware of the membership status of any sample, typically occurring when the defender is tasked with protecting an already trained model without having been involved in its training process. Our approach involves randomly selecting the prediction of one sample from the generated candidates as the output. Although this method does not deliberately restrict the prediction distribution, the reconstruction process inherently contributes to reducing the prediction distribution gap between members and non-members. This effect is achieved as the model encounters newly generated samples during the reconstruction phase, which are distinct from any data it was exposed to during training. This novelty in the samples ensures a more uniform response from the model, diminishing the likelihood of differential predictions between members and non-members.
3.2.3 Prediction Aggregation
@@
By default, we adopt the aforementioned sample selection strategy as it consistently provides privacy protection without altering the model’s accuracy. An alternative option is to aggregate the predictions of all generated samples instead of selecting one. This approach may yield superior results in specific contexts; for example, we observed that direct averaging of predictions from generated samples can enhance both privacy and utility when the diffusion model is trained on sufficient data. Further details on this are provided in Section 4.2.4.
3.2.4 Integration with Other Defenses in a Plug-and-Play Manner
@@
Our method simply employs off-the-shelf diffusion models, offering substantial flexibility in deployment. As outlined in Section 1, we categorize defenses into three deployment phases. Defenses deployed in different stages are compatible and can be used concurrently. Our approach, characterized by minimal defense assumptions, is unique in its deployment at the pre-inference stage and offers plug-and-play capability, allowing for integration with all other existing methods. As illustrated in Figure 2, when combined with training phase defenses, our method can integrated into the inference pipeline immediately after the model has been trained using privacy-preserving techniques.
@@
When combined with post-inference defenses, our approach can be employed to reconstruct input samples within the inference pipeline, followed by the application of the post-inference defenses, such as MemGuard, to introduce noise into the output prediction vector. Our experiments validate the effectiveness of our method in conjunction with other defenses, while also providing new perspectives for the deployment of future defense mechanisms.
4 Evaluation
4.1 Experimental Setup
Datasets. We consider the following benchmark datasets that are widely used in prior works:
@@ 4.1 Experimental Setup
For CIFAR10/CIFAR100, and SVHN datasets, we selected 25,000 samples and 5,000 samples, respectively, to train the target models. The remaining samples from each dataset were reserved as non-members or references used by defenses or attacks.
@@
Models. For target classifiers, we consider the widely used ResNet-18 [21] and DenseNet121 [22] as the target model architecture. In our default setup, each target model are trained for 100 epochs using the Adam optimizer, with a learning rate of 0.001. We apply an L2 weight decay coefficient of 10-6 and use a batch size of 128. As we found the performance of recent works can be affected by the training configurations of the target models, we also employ their settings to demonstrate that our defense enhances their best outcomes.
@@
For the diffusion models integrated into our defense, we train standard DDPMs from scratch using the default hyper-parameters from the original DDPM paper. Unless otherwise mentioned, the diffusion model is trained on the same training data of the associated defended target classifier, following the assumptions that the model onwer has no extra data. Our default configuration involves generating $N=50$ reconstructed images for each sample, with the number of diffusion steps $T$ set to 40. Detailed discussions on the choice of these hyperparameters are provided in Section 4.3.2. To accelerate the inference process, we also employ Denoising Diffusion Implicit Models (DDIM($k$)), where $k$ indicates the inference interval. In our setting, $k$ is set to 10, corresponding to 4 denoising steps when $T=40$.
@@
Attack methods. For evaluation we consider 6 state-of-the-art attack methods. These comprise NN-based attacks [5, 15], four threshold-based attacks (loss, confidence, entropy, M-entropy) and the recent LiRA [29] attack. In line with previous practice, we exclude attacks using only partial model output as they are strictly weaker than attacks above.
@@
Evaluation metrics. We use Top-1 accuracy on validation dataset to quantify model’s utility. For privacy, we default to using five of the described attacks (excluding LiRA) to evaluate two common used metric: (i) attack accuracy and (ii) attack AUC. Additionally, we report the TPR (True Positive Rate) at 0.1% FPR and the TNR (True Negative Rate) at 0.1% FNR on SVHN dataset, using LiRA which is specifically designed to achieve superior results on this metric. We exclusively applied the LiRA attack on the SVHN dataset due to its high computational demands, as it requires the training of over 100 shadow models.
4.2 Experimental Results
@@
Figure 5: Attack accuracy and AUC on three datasets against ResNet under Scenario 1. We report the highest attack accuracy and attack AUC across all attacks. The prediction accuracy delta indicates the prediction accuracy gap compared to the undefended models, with negative numbers indicating a decrease in model utility.

### v2

@@ III-A Intuition Behind Diffence
Figure 3 plots histograms illustrating the differences in predictions for members and non-members on CIFAR-100 without defenses. It can be observed that there are significant disparities in the output distributions between members and non-members, especially in the parametrically modeled confidence, which we refer to as logit in this paper. Attackers can easily exploit these differences to distinguish between members and non-members.
To address these issues, we propose a novel method that leverages input reconstruction to eliminate the prediction disparity between members and non-members. Our core idea is that by reconstructing samples, the model, during classification, encounters samples that are distinct from those it has seen during the training phase, irrespective of whether they are from members or non-members, thereby reducing inconsistencies in predictions. Our proposed method, Diffence, aims to reconstruct the finer details of samples without altering their semantic content. To this end, we employ diffusion models as our generative model, as they align well with our objectives.
Figure 4 provides examples of how Diffence narrows the prediction gap. By reconstructing the input, Diffence ensures that the model’s predictions for both members and non-members become more consistent, thereby mitigating the risk of membership inference attacks. This process reduces the gaps in confidence levels, prediction entropy, and the parametrically modeled confidence.
III-B Our Proposed Diffence
Diffence involves generating multiple reconstructed images and selecting the best one that suits our purpose. Among the generated samples, only those with the same predicted label as the original sample are considered as candidates for selection. This ensures that the model’s predictive outcomes remain unchanged, thereby maintaining the model accuracy. We designed different sample selection strategies for three scenarios based on the assumptions in the defender. These scenarios include a defender having access to both some member and non-member samples, having access only to members, and having access solely to a trained model without any knowledge of samples’ memberships. With more information, defenders can perform a more nuanced analysis of the prediction distribution. This enables a more strategic selection of reconstructed images, forcing the prediction distributions of members and non-members to align more closely, thereby providing robust protection. We provide more details below.
III-B1 Sample Reconstruction
@@
Typically, images display a high response to low-frequency content and a notably weaker response to high-frequency content. This is because of the inherent smoothness of most images. The dominant low-frequency components encapsulate essential visual information, while the high-frequency components, associated with fine details and edges, are comparatively subdued [40].
@@
We denote the diffusion step applied to each sample in Diffence as $T$. By selecting an appropriate value for $T$, we can preserve the essential semantic content of each image while altering its finer details. The impact of the diffusion step $T$ on the performance of Diffence is further explored in Section IV-C2. During the inference phase, the two-step reconstruction procedure reduces the gap between the prediction distributions of members and non-members.
III-B2 Sample Selection
@@
Figure 4: Prediction distribution gap between members and non-members under Diffence across the three scenarios. This figure displays the distribution of predicted logits on the CIFAR-100 dataset using ResNet18 with Diffence in the three scenarios.
We observed that the stochastic nature of sample generation can lead to a decline in sample quality sometimes when only one reconstructed sample is generated for each image, potentially resulting in decreased model test accuracy. To address this issue, we propose generating multiple reconstructed images for each original image. Then in alignment with our defense objective, we carefully select the most appropriate samples to be utilized as the final input sample.
For each original image $I_{i}$ requiring privacy protection, Diffence generates $N$ reconstructed versions, denoted as $R_{i}=\{R_{i1},R_{i2},\ldots,R_{iN}\}$. The model $f$ outputs a prediction for each reconstructed image $R_{ij}$, i.e., $f(R_{ij})$. From these, we identify a subset of reconstructions whose predicted labels match the predicted label of the original image $I_{i}$ forming a set of candidate reconstructions, denoted as $C$. We then select the most appropriate reconstructed image from the candidates. Selecting a sample from this candidate set ensures that the original prediction label of the sample remains unchanged, guaranteeing that Diffence does not impact the model’s accuracy.
@@
Here, $P_{i}$ represents the final prediction for the original image $I_{i}$, derived from the model’s prediction for the selected reconstruction. The selection function, denoted as $\text{select}(\cdot)$, chooses the optimal reconstruction from $R_{i}$ based on a criterion that evaluates the predictions of all $N$ reconstructions. This criterion is aligned with the defense’s objectives and may take into account the defender’s prior knowledge about the prediction distributions of members and non-members. We categorize Diffence into three scenarios based on the information available to the defender and have accordingly designed distinct sample selection strategies for each.
Scenario 1. In this scenario, we assume that the defender has access to subsets of both member and non-member data. The defender initially generates $N$ reconstructed samples for each data point and then plots the prediction distributions for all these reconstructions. Through the analysis of these distribution plots, the defender can identify an optimal interval where the predictions should ideally fall to maximize the reduction of the prediction distribution gap between members and non-members. Specifically, a grid search is conducted over the overlapping regions of member and non-member prediction distribution $[min(logit_{mem}),max(logit_{nonmem})]$ to determine this interval. The chosen interval aims to minimize the Jensen-Shannon (JS) divergence [42] between the prediction distributions of members and non-members within the selected range. This approach is based on the observation that a lower JS divergence between member and non-member prediction distributions typically indicates reduced membership privacy leakage [16]. This correlation has been confirmed by our experiments, as detailed in Section IV-C1.
Defenders require only a small subset of samples to select the optimal interval. After the interval is determined, for each sample requiring protection, the defender randomly selects one sample from the candidate set $C$ that falls within this interval based on the prediction. If none of the candidates are within the interval, the closest one is chosen. Another possible approach involves continuously generating reconstructions until the sample falls within the defined interval. However, we found this method to be inefficient. More details on this are discussed in Section IV-C1.
Figure 4 provides an example of Diffence as applied to the CIFAR-100 dataset. It illustrates the prediction distribution of reconstructed samples for 1000 data points which is used to select the optimal interval. The figure also showcases the distribution of prediction logits for both members and non-members after the selection of the optimal interval using the steps outlined in Diffence. This visualization demonstrates the effectiveness of Diffence in aligning the prediction distributions of member and non-member data, thereby minimizing the potential for privacy leakage.
@@
Scenario 3. In this scenario, the defender is unaware of the membership status of any sample, typically occurring when the defender is tasked with protecting an already trained model without having been involved in its training process. Diffence involves randomly selecting the prediction of one sample from the generated candidates as the output. Although this method does not deliberately restrict the prediction distribution, the reconstruction process inherently contributes to reducing the prediction distribution gap between members and non-members. This effect is achieved as the model encounters newly generated samples during the reconstruction phase, which are distinct from any data it was exposed to during training. This novelty in the samples ensures a more uniform response from the model, diminishing the likelihood of differential predictions between members and non-members.
III-B3 Prediction Aggregation
@@
By default, we adopt the aforementioned sample selection strategy as it consistently provides privacy protection without altering the model’s accuracy. An alternative option is to aggregate the predictions of all generated samples instead of selecting one. This approach may yield superior results in specific contexts; for example, we observed that direct averaging of predictions from generated samples can enhance both membership privacy and model accuracy when the diffusion model is trained on sufficient data. Further details on this are provided in Appendix A-C.
III-B4 Integration with Other Defenses in a Plug-and-Play Manner
@@
Diffence employs off-the-shelf diffusion models, offering substantial flexibility in deployment. As outlined in Section I, we categorize defenses into three deployment phases. Defenses deployed in different stages are compatible and can be used concurrently. Diffence, characterized by minimal defense assumptions, is unique in its deployment at the pre-inference stage and offers plug-and-play capability, allowing for integration with all other existing methods. As illustrated in Figure 2, when combined with training phase defenses, Diffence can integrated into the inference pipeline immediately after the model has been trained using privacy-preserving techniques.
@@
When combined with post-inference defenses, Diffence can be employed to reconstruct input samples within the inference pipeline, followed by the application of the post-inference defenses, such as MemGuard, to introduce noise into the output prediction vector. Our experiments validate the effectiveness of Diffence in conjunction with other defenses, while also providing new perspectives for the deployment of future defense mechanisms.
III-C Defending against Label-Only Attacks
Diffence effectively reduces the prediction distribution gap while preserving the model’s accuracy by not altering the predicted labels for samples. However, this design choice means that Diffence is not directly effective against label-only attacks, which exploit the model’s predicted labels rather than the prediction confidence. Label-only attacks are generally considered weaker and less of a priority in recent defense works. For instance, RelaxLoss [5] excludes label-only attacks from their evaluation, arguing that these attacks are strictly weaker than the ones used to assess model privacy. Similarly, HAMP [6] demonstrates that existing label-only attacks are unsuccessful in a low false positive/negative regime.
Nonetheless, it’s worth noting that previous defenses, such as SELENA [4] and HAMP [6], have shown success in resisting label-only attacks. While Diffence does not directly improve defense against label-only attacks, it can be seamlessly integrated with these existing defenses, enhancing overall protection without sacrificing the model’s utility. Although Diffence does not specifically target label-only attacks, it offers a flexible and non-intrusive approach to enhancing membership privacy that can be effectively combined with other defenses to achieve comprehensive protection.
IV Evaluations
IV-A Experimental Setup
Datasets. We consider three benchmark datasets and two high-resolution datasets :
@@ IV-A Experimental Setup
  • CelebA [25]: CelebA contains over 200,000 high-resolution celebrity images (178x218 pixels), each annotated with 40 attributes. Following prior work [43], we selected 3 attributes to create 8-class labels.
  • UTKFace [26]: A large-scale face dataset with over 20,000 images labeled by age, gender, and ethnicity (resolutions up to 200x200 pixels). We used images from the four largest races (White, Black, Asian, Indian) for race-based labels.
@@
For the CIFAR-10 and CIFAR-100 datasets, we used 25,000 samples each for training the target models. For the SVHN dataset, we used 5,000 samples. For the CelebA and UTKFace datasets, we used 10,000 and 5,000 samples, respectively. The remaining samples from each dataset were reserved as non-members or references used by defenses or attacks.
Models. For target classifiers, we consider the widely used ResNet18 [27] as the target model. Appendix A-A provides experimental results on additional model architectures, including DenseNet121 [28], VGG16 [29], and ViT [30].
@@
In our default setup, each target model is trained for 100 epochs using the Adam optimizer with a learning rate of 0.001. We apply an L2 weight decay coefficient of $10^{-6}$ and use a batch size of 128. For ViT, we use a learning rate of 0.0001. As we found that the performance of some recent works can be affected by the training configurations of the target models, we also employ their training settings to demonstrate that Diffence enhances their best outcomes.
@@
For the diffusion models integrated into Diffence, we train standard DDPMs from scratch using the default hyperparameters from the original DDPM paper [22]. Unless otherwise mentioned, Diffence uses a diffusion model trained on the same training data as the associated defended target classifier, and generates $N=30$ reconstructed images for each sample, with the number of diffusion steps $T$ set to 160. Detailed discussions on the choice of these hyperparameters are provided in Section IV-C2.
In addition to training diffusion models on the same distribution as the target classifier, we also demonstrate the effectiveness of Diffence using publicly available pre-trained diffusion models trained on ImageNet [14]. This illustrates that the diffusion models in Diffence and the target classifier’s training data do not need to be from the same distribution. It also shows that when an off-the-shelf diffusion model for the same dataset is not available, using a pre-trained diffusion model on a different dataset can still provide effective protection. More details are provided in Section IV-B5.
@@
Attack and Defense Methods. For evaluation, we consider six state-of-the-art attack methods: NN-based attacks [1, 17], four threshold-based attacks (loss, confidence, entropy, M-entropy), and the recent LiRA attack [39].
We consider six major defenses: AdvReg [1], MemGuard [2], SELENA [4], RelaxLoss [5], HAMP [6], and DPSGD [3]. For all attack and defense methods, we follow the original papers and setups unless otherwise stated. Detailed information is provided in Appendix A-E.
Following previous practice [5], we exclude label-only attacks in our comparative experiments. Diffence primarily targets attacks beyond label-only attacks, aiming to avoid accuracy loss by not altering predicted labels, and thus does not directly defend against label-only attacks (more discussions are provided in Appendix A-B). However, label-only attacks can be effectively countered by existing defenses such as SELENA [4] and HAMP [6], but defenses against other attacks still need improvement. Our approach is designed to complement these existing defenses on the stronger attacks (i.e., non label-only MIAs), enabling a comprehensive defense strategy that addresses all types of attacks.
@@
Evaluation metrics. For privacy, we default to using five of the described attacks (excluding LiRA) to evaluate two common used metric: (i) attack accuracy and (ii) attack AUC. Additionally, we report the TPR (True Positive Rate) at 0.1% FPR and the TNR (True Negative Rate) at 0.1% FNR on SVHN dataset, using LiRA which is specifically designed to achieve superior results on this metric. We exclusively applied the LiRA attack on the SVHN dataset due to its high computational demands, as it requires the training of over 100 shadow models.
@@
Figure 5: Attack accuracy and AUC on three datasets against ResNet18 under Scenario 1. We report the highest attack accuracy and attack AUC across all attacks. The prediction accuracy delta indicates the prediction accuracy gap compared to the undefended models, with negative numbers indicating a decrease in model accuracy.
TABLE III: Average attack accuracy and AUC across three datasets under three scenarios on ResNet18. The best defense results in each case are highlighted in bold.
Defenses Prediction Accuracy Delta (%) w/o Diffence w/ Diffence (Scenario 1) w/ Diffence (Scenario 2) w/ Diffence (Scenario 3)
Attack AUC (%) Attack Accuracy (%) Attack AUC (%) Attack Accuracy (%) Attack AUC (%) Attack Accuracy (%) Attack AUC (%) Attack Accuracy (%)
Undefended 0 79.14 77.73 68.08 65.41 70.79 67.12 69.12 67.02
SELENA -2.13 62.22 60.92 56.00 55.23 60.30 58.58 57.81 57.16
AdvReg -5.53 61.32 58.94 59.17 57.68 61.33 58.58 60.87 58.62
HAMP -0.23 78.96 76.08 67.60 65.10 71.23 66.61 69.18 66.31
RelaxLoss 0.97 75.81 73.78 67.13 64.75 69.56 66.19 68.60 66.07
DP-SGD -9.13 56.61 56.19 55.47 55.40 58.40 56.92 56.60 56.35
Memguard 0 69.53 68.21 66.76 64.47 67.23 65.30 67.48 65.51
IV-B Experimental Results
IV-B1 Comparison to Baselines
We first conducted an extensive evaluation of Diffence across three benchmark datasets. For each dataset, we evaluated the performance of models across seven different cases including one without any protection and six others, each employing a different defense mechanism. We tested the impact of Diffence on the models’ test accuracy and the changes in attack accuracy and attack AUC under three scenarios introduced in Section 4. Our results show that Diffence consistently enhanced privacy protection across all settings without compromising the model accuracy. We discuss the specific effects of Diffence below.
Figure 5 shows the performance of Diffence against ResNet18 under Scenario 1. We can observe that Diffence consistently reduces the attack’s AUC and accuracy across all cases, particularly against those methods that preserve model usability but offer only limited protection. For instance, when applied to ResNet18, Diffence managed to decrease the attack accuracy for undefended, HAMP, RelaxLoss, and SELENA models by 15.8%, 14.4%, 12.2%, and 9.3%, respectively, on average across three datasets. Similarly, the attack AUCs showed reductions of 14.0%, 14.4%, 11.4%, and 10.0% after employing Diffence. The experimental results demonstrate that Diffence can significantly reduce privacy threats, even with the reliance on only a small amount of additional data (1000 non-members in our experiments). Besides ResNet18, we also tested Diffence on DenseNet121, VGG16, and ViT. The experiments show that Diffence can effectively enhance membership privacy in all these settings (more details in Appendix A-A).
Our experiments suggest that the optimal defense practice is combining Diffence with recent utility-preserving defenses. This combination leverages their advantages in maintaining utility while enhancing privacy protection. For instance, the best privacy-utility trade-offs achieved on VGG16 were from the integrations of Diffence with SELENA and HAMP (Figure 12).
Table III presents the experimental results using ResNet18 under all three different scenarios. We show the average attack AUC and accuracy reduction of Diffence across three datasets in each scenario. The average defense effectiveness in Scenarios 2 and 3 is weaker than in Scenario 1, due to stronger assumptions placed on the defender in these scenarios. However, it is observable that they still achieve significant privacy enhancements across all settings without compromising the model’s accuracy. We observed that some recent defenses, such as HAMP and RelaxLoss, have not achieved satisfactory overall effectiveness. We found their performance heavily depends on the manual selection of hyperparameters. Following their claims about the impact on utility, we selected hyperparameters that matched the test accuracy of undefended models. However, this sometimes resulted in poor defense and even worsened privacy. Diffence can address these privacy issues while maintaining their high utility.
In the research of MIA defense, balancing sufficient privacy protection with model utility has always been one of the most important objectives. Our results suggest that Diffence represents a significant step forward towards the ideal defense.
IV-B2 Evaluation of Confidence Calibration in Model Predictions
Diffence preserves the predicted label of the sample, thereby not affecting the model’s accuracy. Nonetheless, the utility of a model also involves the value of information provided by the output confidence vector. As discussed in Section III-B, Diffence exploits diffusion model to generate reconstructions that closely resemble the original samples, differing only in subtle details (see Figure 1 for examples). We propose that the confidence levels of these reconstructions are reliable indicators and provide a dependable measure of the original sample’s characteristics.
Previous research [32, 35, 44] emphasizes the importance of well-calibrated model outputs, where the predicted class probability should match its actual correctness likelihood. Expected Calibration Error (ECE) [31, 32, 44] is a key metric in this context, quantifying the alignment between predicted probabilities and empirical accuracies. Essentially, ECE measures the average discrepancy between the confidence of a model’s predictions and the true correctness of those predictions, providing a critical assessment of the reliability of the model’s confidence estimates. A lower ECE suggests better alignment of confidence scores with true probabilities, indicating more reliable predictions. ECE is calculated as in equation 6:
$\text{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\left|\text{acc}(B_{m})-\text{conf}(B_{m})\right|$ (6)
where $M$ denotes the number of bins into which the predictions are grouped. $B_{m}$ is the set of samples in the $m$-th bin. $n$ represents the total number of samples. $\text{acc}(B_{m})$ is the accuracy within bin $B_{m}$. $\text{conf}(B_{m})$ is the average predicted confidence for samples in bin $B_{m}$.
Figure 6 shows the ECE of the confidence vectors output by both undefended models and models employing various defenses across three datasets. It is observed that the ECE values of Diffence are very similar to those of the undefended models. For instance, the ECE values for the undefended model and the three Diffence scenarios are 0.139, 0.142, 0.126, and 0.131 on average, respectively. This suggests that Diffence not only maintains the accuracy of the model but also provides meaningful confidence scores.
However, it was observed that some defenses could lead to a significant increase in ECE. For example, compared to the undefended model, the average ECE of the HAMP is 222.3% higher. This could be attributed to HAMP’s strategy of promoting high-entropy confidence vectors as a defense against MIAs, leading to more flattened confidence distributions. Such alterations, while serving a defensive purpose, can yield confidence scores that lack meaningful interpretation, potentially compromising the utility of the model.
In summary, our results validate the effectiveness of Diffence in preserving the utility of confidence scores, and also highlight the importance of evaluating the meaningfulness of confidence scores in the field of MIA defenses.
IV-B3 Effectiveness on Attack TPR and TNR Metrics
Following the methodology of Carlini et al. [39], we evaluate the reliability with which an adversary can compromise the privacy of even a few users in a sensitive dataset, using metrics such as TPR at 0.1% FPR and TNR at 0.1% FNR. We assess the defensive effectiveness of Diffence under the third scenario against these metrics on the SVHN dataset, using six different attacks including LiRA.
As illustrated in Figure 7, Diffence significantly reduces both the attack TPR and TNR in most cases. For example, Diffence decreased the attack TPR under 0.1% FPR and TNR under 0.1% FNR against the undefended model by 63.9% and 56.8%, respectively. When cascaded with other defenses, Diffence reduced the TPR under 0.1% FPR by 52.8% on average across six tested defenses.
Figure 6: Expected Calibration Error (ECE) of various defenses across three datasets. A lower ECE value indicates better-calibrated confidence scores.
Figure 7: Attack TPR and FPR on SVHN dataset. We report the highest attack TPR and attack FPR across all attacks.
Diffence has a smaller impact on methods that already exhibit low TPR; however, these methods often compromise the model’s utility. Consequently, Diffence can be advantageously combined with other defenses that preserve model utility but are relatively weaker in defense effectiveness, to enhance overall protection without significantly sacrificing utility.
IV-B4 Improving Performance Beyond Previous State-of-the-Art